OpenStudy (anonymous):

which equation of a circle with center (5,0) that passes through the point (1,1)

OpenStudy (amistre64):

its really the same as the distance from 5,0 to any (x,y), such that one of them is (1,1) :)

terenzreignz (terenzreignz):

...the only things you really need to determine the equation of a circle is the center and the radius... one of which you already (explicitly) have.

OpenStudy (amistre64):



OpenStudy (amistre64):

\[(x-5)^2+(y-0)^2=r^2\] plug in the point 1,1 to determine r

OpenStudy (anonymous):

(1-5)^2+(1-0)^2=17